OpenStudy (anonymous):

PLEASE HELP!!!! Donna teaches dance in groups of five students. She rents a hall for the dance classes at $450 per month. She charges $40 per hour per student. Donna teaches four groups of students, and her net earnings per month are $4,350. Part A: If each student learns dance for the same number of hours, write an equation that can be used to calculate the number of hours each student learns dance in a month. Part B: How many hours did each student learn dance each month? Show your work and justify each step.

OpenStudy (jdoe0001):

do you know what "net earnings" mean?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

@jdoe0001 ?

OpenStudy (jdoe0001):

hmmm ok... so... net earnings... is what you KEEP after all deductions

OpenStudy (anonymous):

ok, so how does that help me?

OpenStudy (jdoe0001):

so if you make lemonade, and spend $20 in the lemon and the bucket that's $20 that you spend then if at the end of the day, you made $50 then your net is $50 total or "gross earning" minus $20 "expenses" 50 - 20 = 30 <-- net earnings

OpenStudy (anonymous):

ok

OpenStudy (jdoe0001):

well we want to know Donna's expenses, we know her net earnings is 4,350

OpenStudy (anonymous):

right

OpenStudy (anonymous):

so her expenses would be 4350, plus the rent of $450?

OpenStudy (jdoe0001):

well... her expenses are $450 for the hall rent her net earnings are 4,350

OpenStudy (jdoe0001):

so she has 4 groups of students each group is 5 students so 4 * 5 = 20 students total each hour of class is $40 per student now we dunno how many hours they learn dance for so let's say they learn dance "x" hours

OpenStudy (anonymous):

ok

OpenStudy (jdoe0001):

so 20 students $40 per hour, so for "x" hours that'll be 40x her expenses are 4,350 so one can say (students) * ( 40 * hours) - (hall rent ) = 4,350

OpenStudy (jdoe0001):

or \(\bf 20\cdot 40x-450=4,350\implies 800x-450=4,350\)

OpenStudy (jdoe0001):

the 450 is subtracted because she's not earning 450, she's spending 450 to rent the dance hall

OpenStudy (jdoe0001):

and to get part B "How many hours" just solve for "x"

OpenStudy (anonymous):

THANKS YOU SO MUCH!!!!! i had 2 eat my lunch, so I was away from my computer ;)

OpenStudy (jdoe0001):

yw